# Fast-weight Product Key Memory

### v1

@@ Abstract
Sequence modeling layers in modern language models typically face a trade-off between storage capacity and computational efficiency. While Softmax attention offers unbounded storage at prohibitive quadratic costs, linear variants provide efficiency but suffer from limited, fixed-size storage. We propose Fast-weight Product Key Memory (FwPKM), a novel architecture that resolves this tension by transforming the sparse Product Key Memory (PKM) from a static module into a dynamic, “fast-weight” episodic memory. Unlike PKM, FwPKM updates its parameters dynamically at both training and inference time via local chunk-level gradient descent, allowing the model to rapidly memorize and retrieve new key-value pairs from input sequences. Experiments reveal that FwPKM functions as an effective episodic memory that complements the semantic memory of standard modules, yielding significant perplexity reductions on long-context datasets. Notably, in Needle in a Haystack evaluations, FwPKM generalizes to 128K-token contexts despite being trained on only 4K-token sequences.
Abstract
Sequence modeling layers in modern language models typically face a trade-off between storage capacity and computational efficiency. While Softmax attention offers unbounded storage at prohibitive quadratic costs, linear variants provide efficiency but suffer from limited, fixed-size storage. We propose Fast-weight Product Key Memory (FwPKM), a novel architecture that resolves this tension by transforming the sparse Product Key Memory (PKM) from a static module into a dynamic, “fast-weight” episodic memory. Unlike PKM, FwPKM updates its parameters dynamically at both training and inference time via local chunk-level gradient descent, allowing the model to rapidly memorize and retrieve new key-value pairs from input sequences. Experiments reveal that FwPKM functions as an effective episodic memory that complements the semantic memory of standard modules, yielding significant perplexity reductions on long-context datasets. Notably, in Needle in a Haystack evaluations, FwPKM generalizes to 128K-token contexts despite being trained on only 4K-token sequences.
@@ 1 Introduction
Sequence modeling layers, or token mixers, are the foundational components in modern language models. The most successful architectures today can be fundamentally understood as forms of associative memory (Transformer; ParallelDeltaNet; GDN; RWKV7; Mamba2), characterized by their ability to maintain key-value associations, execute retrieval, and perform memorization (KVMemInBrain).
Within this framework, existing layers lie on a spectrum defined by the trade-off between storage capacity and computational efficiency. Standard softmax attention (Transformer) acts as an associative memory with unbounded storage, yet its computational cost becomes increasingly prohibitive as the sequence length grows (TransformerAssocMem). Conversely, linear attention variants (LinearAttention; DeltaNetLSTM; Mamba; Mamba2; GDN; TTT; Titans) provide efficient, sub-quadratic mechanisms but rely on fixed storage capacities that often struggle to capture the same depth of information.
We focus our investigation on resolving this specific tension: balancing large-scale storage with low computational overhead. We posit that an ideal associative memory should satisfy four key properties:
  1. nosep
    Key-value Association: The ability to link keys to values.
  2. nosep
    Large Storage: Capacity that is extensive, if not unbounded.
  3. nosep
    Low Cost: Sub-quadratic computational complexity w.r.t. input length.
  4. nosep
    Retrieval and Memorization: The capability to retrieve information and, crucially, memorize new key-value pairs from inputs at any time.
Product Key Memory (PKM, PKM) is an architecture that elegantly satisfies the first three properties. Its sparse key-value design handles an enormous number of memory slots (e.g., $N=10^{6}$) with fixed and low computation. However, PKM was originally designed as a “slow-weight” channel mixer – similar to Feed-Forward Networks (FFN) – meaning it is updated only during training and remains frozen during inference. Consequently, it lacks the ability to rapidly adapt to new inputs during deployment, failing property 4.
In this paper, we propose to convert PKM from a static, slow-weight module into Fast-weight Product Key Memory (FwPKM). By redesigning PKM to update its parameters dynamically at both training and inference time, we enable it to function as a high-fidelity episodic memory. FwPKMcan store “episodes” directly from input sequences and carry that memory across different contexts, offering a promising new path for continual learning and personalized AI agents.
Table 1: Comparison of PKM and FwPKM
PKM FwPKM
Weight Type Slow Weights Fast Weights
Similar Modules FFN, MoE-FFN, PKM Softmax Attn., Linear Attn., DeltaNet
Role Channel Mixer. Mixes features within a single token representation. Token Mixer. Mixes information across time steps (sequence positions).
Parameter Update Updated at training; Frozen at inference. Updated at both training and inference.
Memory Horizon Long-term (Semantic). Stores dataset-wide facts and general rules (e.g., world knowledge). Short-term (Episodic). Stores context-specific variable bindings (the Context Window).
Learning Objective Global Objective of next token prediction Local Objective of memorization
2 Product Key Memory
Top-$k$ Key-value Memory
A standard key-value memory consists of a key matrix $K\in\mathcal{R}^{N\times D_{K}}$ and a value matrix $V\in\mathcal{R}^{N\times D_{V}}$, where $N$ represents the number of memory slots and $D_{\{K,V\}}$ are the hidden dimensions. A common approach to learning a large memory without sacrificing computation efficiency is to exploit sparsity via a Top-$k$ operation (MemNet-JW; SparseMemRae). Given an input query vector $\mathbf{q}$, the model computes a score $s_{i}$ for each memory slot as the inner product between the query and the keys. A Top-$k$ operation $\mathcal{T}_{k}$ then selects the indices of the $k$ slots with the highest scores. The selected scores are normalized via softmax to produce weights $\{s^{\prime}_{i}\}$, and the retrieval output $\hat{\mathbf{v}}$ is computed as the weighted sum of the corresponding value slots.
$s_{i}$ $=\mathbf{q}^{\top}K_{i}$ (1)
$\mathcal{I}$ $=\mathcal{T}_{k}(\mathbf{s})$ (2)
$\{s^{\prime}_{i}\}$ $=\mathtt{softmax}(\{s_{j}\}_{j\in\mathcal{I}})$ (3)
$\hat{\mathbf{v}}$ $=\textstyle\sum_{i\in\mathcal{I}}V_{i}s^{\prime}_{i}$ (4)
Product Key Memory
While the Top-$k$ operation restricts the number of accessed memory slots, it still requires computing scores for all $N$ key rows to find the top candidates. This $O(N)$ complexity prohibits scaling to enormous memory sizes (e.g. $N=10^{6}$). PKM proposed Product Key Memory (PKM) to address this problem. In PKM, the query vector $\mathbf{q}$ is decomposed into two sub-queries $\mathbf{q}^{1}$ and $\mathbf{q}^{2}$ as shown below in Equation 5, where $\cdot$$\cdot$] is concatenation. Instead of a single large key matrix, PKM maintains two smaller sub-key matrices $K^{\{1,2\}}$ each of size $\sqrt{N}\times D_{K}$. The memory slots are arranged in a Cartesian grid of size $\sqrt{N}\times\sqrt{N}$, and the slot at index $(i,j)$ corresponds to the interaction between the $i$-th sub-key from $K^{1}$ and the $j$-th sub-key from $K^{2}$. The score for this slot is defined as the sum of the scores of its corresponding sub-queries: $s_{(i,j)}=s_{i}^{1}+s_{j}^{2}$. Importantly, under this formulation, the Top-$k$ elements of this Cartesian product can be obtained without computing all $N$ scores. Instead, it first identifies the Top-$k$ indices for each sub-query independently and then searches within the smaller length-$k^{2}$ Cartesian product of these selected sets.
$[\mathbf{q}^{1};\mathbf{q}^{2}]$ $=\mathbf{q}$ (5)
$s^{\{1,2\}}_{i}$ $=\mathbf{q}^{\{1,2\}\top}K^{\{1,2\}}_{i}$ (6)
$\mathcal{I}^{\{1,2\}}$ $=\mathcal{T}_{k}(\mathbf{s}^{\{1,2\}})$ (7)
$\mathcal{I}$ $=\mathcal{T}_{k}(\{s_{i}^{1}+s_{j}^{2}|i\in\mathcal{I}^{1},j\in\mathcal{I}^{2}\})$ (8)
Once the final indices $\mathcal{I}$ are selected, the remaining retrieval process (i.e. score normalization and weighted sum) keeps the same.
$\hat{\mathbf{v}}$ $=\mathtt{PKM}(\mathbf{q};K,V)=\textstyle\sum_{i\in\mathcal{I}}V_{i}s^{\prime}_{i}$ (9)
The PKM architecture provides an elegant way to build a large-scale memory module at a low cost. A $10^{6}$-slot memory can be indexed using two small key matrices of size $10^{3}$, requiring only $2\times 10^{3}$ score calculations. Furthermore, PKM naturally maintains a mapping between keys and values, thus making it an ideal foundation for the high-capacity associative memory we want to build.

### v2

@@ Abstract
Sequence modeling layers in modern language models typically face a trade-off between storage capacity and computational efficiency. While softmax attention offers unbounded storage at prohibitive quadratic cost, linear variants are more efficient but suffer from limited, fixed-size storage. We introduce Fast-weight Product Key Memory (FwPKM), a sparse fast-weight memory layer that resolves this tension. FwPKM updates sparsely activated parameters at both training and inference time using chunk-level gradient descent on a local memory-rewrite objective. This performs Test-Time Training (TTT)-style gradient updates on activated slots in a sparse memory, enabling rapid memorization and retrieval of many new key–value associations while keeping per-token compute low and fixed. Experiments show that FwPKM functions as an effective episodic memory that complements the semantic memory of standard modules, yielding significant perplexity reductions on long-context datasets. Notably, in Needle-in-a-Haystack evaluations, FwPKM generalizes to 128K-token contexts despite being trained on only 4K-token sequences.
@@ 1 Introduction
Long-context language modeling often requires remembering specific, local facts (e.g. a name, a variable binding, a constraint) introduced tens of thousands of tokens earlier, and retrieving them when needed. The core sequence modeling layers, or token mixers, can be viewed as associative memory systems (Transformer; ParallelDeltaNet; GDN; RWKV7; Mamba; Mamba2; TTT; Titans). In this framework, information from past tokens is encoded into storage (i.e. the process of memorization) and later integrated into a prediction step by querying that storage (i.e. the process of retrieval).
The current landscape is defined by a trade-off between memory capacity and computational efficiency. Softmax self-attention (Transformer; TransformerAssocMem) memorizes by explicitly storing a key-value pair for every past token and retrieves by comparing the current query to all stored keys, producing a weighted sum of values. This yields high-fidelity retrieval but at quadratic cost with sequence length. In contrast, modern Recurrent Neural Networks (RNNs) such as linear attention (LinearAttention), Mamba2 (Mamba2), and DeltaNet (FWLinearAttention; ParallelDeltaNet) compress history into a fixed-size state using learned update rules (e.g., Hebbian-style accumulation or Delta-rule error correction) and retrieve via a state readout in the form of vector–matrix multiplication. While this enables efficient, sub-quadratic retrieval, the fixed state capacity inherently limits the depth of information the model can retain compared to softmax attention.
To address the capacity limitations of these modern RNNs, recent work on Test-Time Training (TTT, TTT; TTTPlus) and Titans (Titans) replaces the fixed state matrix with a “fast-weight” (FW-Hinton; FW-Schmidhuber) neural network (typically an MLP). Similar to a state matrix, this MLP adapts to input sequences on-the-fly by minimizing a reconstruction loss. However, these approaches face a scaling bottleneck: dense network structure. An MLP’s computation scales linearly with its parameter count. To store the massive amount of information required for long-context tasks, the fast-weight MLP must be large; yet, updating and querying a large, dense MLP frequently is computationally infeasible.
In this paper, we show that sparse associative memory offers a way around this constraint. Sparse memory architectures such as Product Key Memory (PKM, PKM) can maintain a very large number of memory parameters while activating only a small subset per token. We introduce Fast-weight Product Key Memory (FwPKM), which transforms PKM from a static, slow-weight module into a dynamic fast-weight episodic memory. By redesigning the sparse memory to update its key/value matrices via gradient descent at inference time, we enable FwPKM to memorize vast amounts of context with fixed per-token compute. This distinction allows us to conceptually separate memory roles: standard slow weights serve as Semantic Memory (storing dataset-wide facts), while FwPKM serves as Episodic Memory (storing context-specific bindings).
Our main contributions are threefold.
  • Architecture. We introduce Fast-weight Product Key Memory (FwPKM), a new sequence modeling layer that enables “Sparse Test-Time Training” – TTT-style gradient updates over a sparse memory. We detail the architectural modifications and optimization objectives required to make sparse fast weights stable and effective (Section 3).
  • Performance. We demonstrate that FwPKM acts as a robust episodic memory when plugged into hybrid backbones. In Needle-in-a-Haystack evaluations, FwPKM generalizes to 128K-token contexts despite being trained on only 4K sequences, with iterative reading boosting retrieval accuracy from $10\%$ to $70\%$ (Section 4.3). FwPKM also consistently reduces perplexity on long-context datasets like LC64 and Longbench (Section 4.2, 4.4). Furthermore, we analyze continual learning capabilities, finding that while the model adapts to new domains online, it faces challenges with retention, motivating future work in memory consolidation (Section 4.5).
  • Analysis. We provide interpretability analyses showing that FwPKM explicitly captures informative entities and novel patterns (Section 5), alongside a cost analysis highlighting where sparse memory helps and where faster kernels are needed for broader scaling (Section 6).
2 Preliminary
We briefly review sparse key-value memories and Product Key Memory (PKM, PKM), which we later extend into an online-updated fast-weight module.
2.1 Sparse Top-$k$ key-value memory
A standard key-value memory consists of a key matrix $K\in\mathbb{R}^{N\times d_{k}}$ and a value matrix $V\in\mathbb{R}^{N\times d_{v}}$, where $N$ is the number of memory slots and $d_{\{k,v\}}$ are the hidden dimensions. Given a query vector $\mathbf{q}\in\mathbb{R}^{d_{k}}$, the model scores each slot (e.g., via dot product),
$s_{i}$ $=\mathbf{q}^{\top}K_{i},$ (1)
selects the indices of the Top-$k$ scores $\mathcal{I}$,
$\mathcal{I}$ $=\texttt{Top-}k(\{s_{i}\}_{i=1}^{N}).$ (2)
Then it computes normalized scores $\{s^{\prime}_{i}\}$ via softmax,
$\{s^{\prime}_{i}\}$ $=\mathtt{softmax}(\{s_{j}\}_{j\in\mathcal{I}}),$ (3)
and retrieves a weighted sum of values,
$\hat{\mathbf{v}}$ $=\textstyle\sum_{i\in\mathcal{I}}s^{\prime}_{i}V_{i}.$ (4)
Top-$k$ sparsity reduces the number of accessed value rows, but still requires scoring all $N$ keys to identify the top candidates. This linear $O(N)$ complexity prohibits scaling to massive memory sizes (e.g. $N\approx 10^{6}$).
2.2 Product Key Memory (PKM)
Product Key Memory (PKM) reduces the scoring cost by factorizing the keys into two smaller codebooks. PKM splits the query into two sub-queries $\mathbf{q}=[\mathbf{q}^{(1)};\mathbf{q}^{(2)}]$, where $\mathbf{q}^{(1)},\mathbf{q}^{(2)}\in\mathbb{R}^{d_{k}/2}$, and maintains two sub-key matrices $K^{(1)},K^{(2)}\in\mathbb{R}^{\sqrt{N}\times d_{k}/2}$. It scores each codebook independently,
$s^{(1)}_{i}$ $=\mathbf{q}^{(1)\top}K^{(1)}_{i},\quad s^{(2)}_{j}=\mathbf{q}^{(2)\top}K^{(2)}_{j}$, (5)
selects Top-$k$ indices for each codebook,
$\mathcal{I}^{(1)}=\texttt{Top-}k(\{s^{(1)}_{i}\}_{i=1}^{\sqrt{N}}),\quad\mathcal{I}^{(2)}=\texttt{Top-}k(\{s^{(2)}_{i}\}_{i=1}^{\sqrt{N}}),$ (6)
and then searches over the restricted Cartesian product $\mathcal{I}^{(1)}\times\mathcal{I}^{(2)}$ using additive pairwise scores
$s_{i,j}$ $=s^{(1)}_{i}+s^{(2)}_{j},\quad(i,j)\in\mathcal{I}^{(1)}\times\mathcal{I}^{(2)}.$ (7)
Finally, PKM selects the Top-$k$ pairs from this restricted set, applies softmax over the selected pair scores, and retrieves from the corresponding rows of the value table $V\in\mathbb{R}^{N\times d_{v}}$, where 1-indexed pair $(i,j)$ maps to row $((i-1)\times\sqrt{N}+j)$.
This decomposition of key space yields an efficient Top-$k$ retrieval that avoids scoring all $N$ slots. It allows PKM to index millions of slots (e.g., $10^{6}$) using only thousands of scores ($2\times 10^{3}$), making it an ideal choice for high-capacity associative memory.